Sam is four years younger than Jo. But in five years time Jo will be twice Sam age now. How old are they now?
Answer by ikleyn(52790)


Let  x  be the Sam's age now.
Then Jo's age now is  x+4,  according to the condition.

In  5  years Jo's age will be  (x+4)+5,  and it is twice  Sam's age now:

(x+4)+5 = 2x.

It gives

x + 9 = 2x,
x = 9.

So,  Sam is  9  years old now.  Jo is  9 + 4 = 13  years old now.

Answer.  Sam is  9  years old now.  Jo  13  years old now.